### v1

@@ 6.1 Comparing User-Proxy LLMs on Human-Likeness
Figure 3 compares five user-proxy LLMs across four datasets with a fixed judge (Claude-4-Sonnet) and assistant (GPT-4o). The left three columns of subplots report judge-based realism (GTEval, PI $\Delta w$, RNR; higher is better); the right three columns report human-anchored lexical diversity (MATTR, HD-D, Yule’s K; best at the human z-score = 0). Dashed red (HH) and green (PP) lines provide calibration controls; error bars denote 95% CIs.
@@
Judge realism is consistent across datasets. Across GTEval, PI $\Delta w$, and RNR, Gemini-2.5-Pro and Claude-4-Sonnet are the most human-like on every dataset, with GPT-4o competitive but generally behind, and GPT-OSS-120B and GPT-5 trailing. On ClariQ and QULAC, Claude-4-Sonnet/Gemini-2.5-Pro approach the HH ceiling under RNR, while PI $\Delta w$ shows clear positive win margins for these models and negative/near-zero deltas for the rest. The agreement among the three judge metrics indicates a stable ordering not driven by any single rubric.
Figure 4: Judge sensitivity of judge-realism metrics on ChatbotArena with assistant & user-proxy both set to GPT-4o; bars vary the judge model. Error bars are 95% CIs.
@@
Diversity shows strong dataset effects and a realism–diversity tension. Lexical diversity diverges from the realism ranking. On ClariQ, most notably Claude-4-Sonnet and GPT-5 exceed the human anchor on MATTR/HD-D and exhibit lower Yule’s K, indicating a more diverse vocabulary than human questioners in this information-seeking regime. In contrast, QULAC exhibits a uniform diversity deficit: all proxies fall below the human baseline on MATTR and HD-D (with positive shifts in Yule’s K), suggesting more templated clarifications than humans. ChatbotArena and OASST1 sit between these extremes with smaller deviations around the human baseline. Overall, Gemini-2.5-Pro yields the strongest diversity alignment with humans across datasets; Claude-4-Sonnet generally tends to overshoot the human baseline, whereas GPT-4o is better calibrated to stay closer to the human anchor with fewer extremes. GPT-5 and GPT-OSS-120B have generally larger diversity difference compared to human baselines.
@@
Judge realism and diversity are partially decoupled. High judge realism does not guarantee human-level diversity. Claude-4-Sonnet and Gemini-2.5-Pro lead on GTEval/PI/RNR but under-shoot diversity on QULAC, indicating judges favor intent/style over surface variety. Conversely, GPT-4o shows more stable diversity with moderate judge-realism gains, underscoring that diversity alone is insufficient to achieve judge indistinguishability.
@@
Uncertainty and robustness. Confidence intervals are generally narrow. When overlaps occur (e.g., GTEval on ChatbotArena), rank differences are small and concur with PI/RNR trends. For lexical-diversity panels, confidence intervals occasionally widen, notably for Yule’s K, as length sensitivity inflates variance. Overall, Figure 3 supports three takeaways: (i) Gemini-2.5-Pro and Claude-4-Sonnet are reliably the most human-like by judge criteria; (ii) diversity depends strongly on dataset regime and proxies often lag on clarification-centric tasks like QULAC; and (iii) realism and diversity capture complementary facets of user-proxy quality, motivating the use of both families of metrics.
Figure 5: Judge–human correlation on ChatbotArena: Correlation of Claude-4-Sonnet judge scores & human scores for GTEval and PI, evaluated on Gemini-2.5-Pro user-proxy outputs (N=100 per metric). Solid line: linear fit; dashed: identity. Point size encodes local sample density. All correlations p << 0.001.
6.2 Judge Sensitivity Analysis
@@
With the assistant and user-proxy fixed to GPT-4o, Figure 4 shows substantial judge sensitivity on realism scores. On GTEval, scores spread widely ($\approx\!0.45\text{--}0.81$), with GPT-4o as judge yielding the highest value. PI is the most volatile: Gemini-2.5-Pro & Claude-4-Sonnet produce near-zero or negative win deltas, while GPT-5 & GPT-4o are clearly positive, suggesting family/self-preference or rubric alignment effects (auxiliary evaluations in Appendix C). RNR saturates near the ceiling for GPT-4o & GPT-5 judges ($\approx\!0.96\text{--}0.98$) and is lower for others ($\approx\!0.79\text{--}0.82$), indicating judge-dependent sensitivity: PI >> GTEval >> RNR.
@@
These differences imply that conclusions drawn from a single judge can shift both the absolute level and the ordering of models. In practice, we recommend evaluating with multiple judges and applying HH/PP calibration when feasible, or normalizing per-judge before aggregation, to avoid over-interpreting judge-specific biases. Empirically, Claude-4-Sonnet behaves as a conservative yet stable judge, yielding non-saturated, well-separated scores across metrics, which is why we use Claude-4-Sonnet as a judge in the results reported in Figure 3.
6.3 Human-Judge Correlation
@@
We validate judge reliability by correlating Claude-4-Sonnet judge scores with blinded human expert annotations on ChatbotArena. We stratify 100 episodes per metric (GTEval, PI) by judge score and conversation length, then compute Spearman’s $\rho$ Kokoska and Zwillinger (2000), Pearson’s $r$ Kowalski (2018), and Kendall’s $\tau$ Kendall (1938). As shown in Figure 5, GTEval shows strong alignment with humans, while PI exhibits a moderate correlation, consistent with the added difficulty of pairwise judgments. Overall, these results indicate that judge scores correlate with human perceptions of user-proxy quality.
6.4 Telemetry, Scaling, and Cost
@@
Figures 6, 7, and 8 summarize the practical side of evaluation. Per-episode telemetry (Fig. 6) shows that token usage is dominated by the judge, with the user-proxy & assistant contributing a smaller but non-negligible share; datasets differ markedly, OASST1 drives the highest token counts while ClariQ yields the largest end-to-end latency.
@@
Figure 6: Avg per-episode telemetry for GPT-4o as user-proxy & assistant, and judged by Claude-4-Sonnet across 4 datasets on 6 metrics (Fig. 3); only non-cached episodes considered. Left: Token usage by role (darker: input, lighter: output). Right: Cumulative latency per episode.
Figure 7: Throughput vs. concurrency for GTEval on ChatbotArena (async backend, cache off); only the judge varies. User-proxy and assistant fixed to GPT-4o.
@@
Figure 8: Cost-quality trade-off for PI: cost per evaluation (USD) vs. PI $\Delta w$ ($\uparrow$ better). Judge = Claude-4-Sonnet; assistant = GPT-4o; temperature = 0; cache off. Markers denote user-proxies; labels denote datasets. Dashed line: Pareto frontier. See Table 4 for model pricing.
@@
Scaling on an async backend (cache off) reveals clear judge-dependent throughput (Fig. 7): GPT-4o as judge attains the highest episodes/min and continues to benefit up to high concurrency, Claude-4-Sonnet scales steadily to mid/high throughput, and Gemini-2.5-Pro plateaus earlier. The cost-quality frontier (Fig. 8) clarifies trade-offs for PI evaluation: User-proxies Gemini-2.5-Pro and Claude-4-Sonnet offer attractive Pareto points (good PI $\Delta w$ at moderate cost), and GPT-5 generally incurs higher cost with weaker PI gain. Overall, user-proxy Gemini-2.5-Pro, along with Claude-4-Sonnet as judge, provides a balanced choice for large runs, where one maximizes throughput and another prioritizes peak PI quality.
7 Conclusion & Future Work
MirrorBench introduces a principled, systems-oriented framework for evaluating the human-likeness of user-proxy agents, explicitly decoupled from downstream task success. The framework provides a modular six-layer stack with typed interfaces, metadata-aware registries, a compatibility-checking planner, multi-backend execution, caching, and observability; it supports pluggable proxies, datasets, tasks, and metrics, producing variance-aware, reproducible results at scale. Empirically, we report (GTEval, PI, RNR) alongside human-anchored lexical diversity (MATTR, HD-D, Yule’s K), revealing a realism-diversity tension across domains, and show that absolute scores and model orderings can shift with the choice of judge, underscoring the need for calibration (HH/PP) and multi-judge reporting. Finally, telemetry (tokens, latency) and cost analyses make the practical trade-offs of large-scale evaluation explicit.
@@ 7 Conclusion & Future Work
Limitations. Our results rely on LLM-as-judge metrics that can exhibit model-family bias; HH/PP controls help, but residual bias and prompt sensitivity remain. Our experiments use a single seed; assistant models are largely fixed; and coverage is limited to four English-centric datasets. Finally, diversity metrics capture surface variation and do not fully reflect discourse phenomena (e.g., repair, initiative).
Future Work. (i) Judges: multi-judge ensembles, stronger calibration/normalization, and significance across runs; (ii) Metrics: discourse and interaction-level realism (turn-taking, self-correction, persistence), robustness to prompt/goal perturbations; (iii) Datasets: broader domains (customer support, safety-critical, adversarial contexts) and multilingual corpora; (iv) Systems: distributed backends and first-class report generation; (v) Reproducibility & openness: a live-updating benchmark for the research community.
We aim for MirrorBench to serve as a practical evaluation harness and standardized benchmark for measuring user-proxy realism.
Appendix A Layer-by-Layer Architecture Details
As summarized in Section 3.1, the framework is designed in 6-layers starting from low-level execution backends & persistence to the high-level abstract API interface and CLI. Each of the layers in the architecture shown in Figure 1 are built on top of layers beneath it. Few modules in the mid-upper layers in the architecture also utilizes their sibling modules on the same layer. Here, we briefly describe each layer and the modules built as part of them.
A.1 Execution Backends & Persistence
@@
Execution Backend: It is responsible for taking decomposed units of tasks as input and send them for execution on the respective execution engine. The framework provides abstractions to support multiple execution backends to execute evaluation jobs at scale. The abstract interface provides uniform schema for implementations to extend the support for new execution engines. The framework supports fully-functional synchronous & asynchronous execution backends, and provides stub implementation to extend for Ray backend.
Persistence: All evaluation artifacts are persisted in a SQLite database organized around a three-level hierarchy: runs define experiment configurations, units specify individual evaluation tasks (parameterized by user-proxy, dataset, metric, and seed), and episodes capture per-trial results including execution duration, metric values, and telemetry. Aggregate statistics (mean, standard deviation, confidence intervals) are materialized in a separate metrics table to support efficient querying, while composite scores across multiple metrics are stored in scorecards for holistic comparisons. The complete schema is detailed in Table 1.
Table 1: Complete MirrorBench Run Database Schema
@@ A.2 Core Engine
Data Models: All the data structures for messages, episodes, evaluation units and run manifests are defined with typed data models using TypedDict and pydantic. We formally explain the context of these different data objects later.
Registry Builder: The framework provides certain high-level components which requires module-specific registries. This layer provides the abstract registry builder that supports creating metadata-aware registry for the downstream modules. Such registries can avoid silent errors since we can ensure certain constraints on the classes being registered with the metadata.
Configuration & Manifest Management: A typical evaluation job requires configuring user-proxy, datasets, metrics, scorecards etc. We define standard configuration schema for each of such components. Moreover, upon invocation of job or during dryrun, the system decomposes evaluation task into small evaluation units. In order to support parallelization (for async and remote backends), we persist self-contained information about the executable units as manifest files.
A.3 Orchestration

### v2

@@ 4.1. Comparison of User-Proxy LLMs
Figure 2 compares five user-proxy LLMs across four datasets with a fixed judge (Claude-4-Sonnet) and assistant (GPT-4o). The left three columns of subplots report judge-based realism (GTEval, PI $\Delta w$, RNR; higher is better); the right three columns report human-anchored lexical diversity (MATTR, HD-D, Yule’s K; best at the human z-score = 0). Dashed red (HH) and green (PP) lines provide calibration controls; error bars denote 95% CIs. We provide the assistant the reference chat history to anchor the rollout to the same trajectory as the original conversation; while LLM hallucinations can occasionally induce drift, a manual inspection found such divergence in mere $1\%$ of samples evaluated as part of 2.
Judge realism is consistent across datasets
@@
Across GTEval, PI $\Delta w$, and RNR, Gemini-2.5-Pro and Claude-4-Sonnet are the most human-like on every dataset, with GPT-4o competitive but generally behind, and GPT-OSS-120B and GPT-5 trailing. On ClariQ and QULAC, Claude-4-Sonnet/Gemini-2.5-Pro approach the HH ceiling under RNR, while PI $\Delta w$ shows clear positive win margins for these models and negative/near-zero deltas for the rest. The agreement among the three judge metrics indicates a stable ordering not driven by any single rubric.
Diversity shows strong dataset effects and a realism–diversity tension
@@
Lexical diversity diverges from the realism ranking. On ClariQ, most notably Claude-4-Sonnet and GPT-5 exceed the human anchor on MATTR/HD-D and exhibit lower Yule’s K, indicating a more diverse vocabulary than human questioners in this information-seeking regime. In contrast, QULAC exhibits a uniform diversity deficit: all proxies fall below the human baseline on MATTR and HD-D (with positive shifts in Yule’s K), suggesting more templated clarifications than humans. ChatbotArena and OASST1 sit between these extremes with smaller deviations around the human baseline. Overall, Gemini-2.5-Pro & GPT-4o yield the strongest diversity alignment with humans across datasets, whereas Claude-4-Sonnet generally tends to overshoot the human baseline. GPT-5 and GPT-OSS-120B have generally larger diversity difference compared to human baselines.
Judge realism and diversity are partially decoupled
@@
High judge realism does not guarantee human-level diversity. Claude-4-Sonnet and Gemini-2.5-Pro lead on GTEval/PI/RNR but under-shoot diversity on QULAC, indicating judges favor intent/style over surface variety. Conversely, GPT-4o shows more stable diversity with moderate judge-realism gains, underscoring that diversity alone is insufficient to achieve judge indistinguishability.
Uncertainty and robustness.
@@
Confidence intervals are generally narrow. When overlaps occur (e.g., GTEval on ChatbotArena), rank differences are small and concur with PI/RNR trends. For lexical-diversity panels, confidence intervals occasionally widen, notably for Yule’s K, due to very short the user-side text (e.g., due to few turns) that increases estimator variance. Overall, Figure 2 supports three takeaways: (i) Gemini-2.5-Pro and Claude-4-Sonnet are reliably the most human-like by judge criteria; (ii) diversity depends strongly on dataset regime and proxies often lag on clarification-centric tasks like QULAC; and (iii) realism and diversity capture complementary facets of user-proxy quality, motivating the use of both families of metrics.
4.2. Judge Sensitivity Analysis
@@
With the assistant and user-proxy fixed to GPT-4o, Figure 3 shows substantial judge sensitivity on realism scores. On GTEval, scores spread widely ($\approx\!0.45\text{--}0.81$), with GPT-4o as judge yielding the highest value. PI is the most volatile: Gemini-2.5-Pro & Claude-4-Sonnet produce near-zero or negative win deltas, while GPT-5 & GPT-4o are clearly positive, suggesting family/self-preference or rubric alignment effects. RNR saturates near the ceiling for GPT-4o & GPT-5 judges ($\approx\!0.96\text{--}0.98$) and is lower for others ($\approx\!0.79\text{--}0.82$), indicating judge-dependent sensitivity: PI >> GTEval >> RNR.
Figure 3. Judge sensitivity of judge-realism metrics on ChatbotArena with assistant & user-proxy both set to GPT-4o; bars vary the judge model. Error bars are 95% CIs.
@@
These differences imply that conclusions drawn from a single judge can shift both the absolute level and the ordering of models. Since LLM judges can exhibit systematic biases (Thakur et al., 2025; Sahoo et al., 2025), they may behave either overly generous or overly conservative. Thus in practice, we recommend evaluating with multiple judges and applying HH/PP calibration when feasible, or normalizing per-judge before aggregation, to avoid over-interpreting judge-specific biases. Empirically, Claude-4-Sonnet behaves as a conservative yet stable judge, yielding non-saturated, well-separated scores across metrics, which is why we use Claude-4-Sonnet as a judge in the results reported in Figure 2.
4.3. Human-Judge Correlation
Although judge-sensitivity analysis (§4.2) helps characterize a judge’s behavior on our task, it can still be difficult to interpret, since whether a judge appears “generous” or “conservative” is itself task-dependent. We therefore additionally assess judge reliability via human-judge correlation.
Figure 4. Judge–human correlation on ChatbotArena: Correlation of Claude-4-Sonnet judge scores & human scores for GTEval and PI, evaluated on Gemini-2.5-Pro user-proxy outputs (N=100 per metric). Solid line: linear fit; dashed: identity. Point size encodes local sample density. All correlations p << 0.001.
@@ 4.3. Human-Judge Correlation
We validate judge reliability by correlating Claude-4-Sonnet judge scores with blinded human expert annotations on ChatbotArena. We stratify 100 episodes per metric (GTEval, PI) by judge score and conversation length, then compute Spearman’s $\rho$ (Kokoska and Zwillinger, 2000), Pearson’s $r$ (Kowalski, 2018), and Kendall’s $\tau$ (Kendall, 1938). As shown in Figure 4, GTEval aligns strongly with human judgments, while PI exhibits a moderate correlation. Note that PI is inherently harder to annotate because it requires pairwise comparisons; thus, a moderate correlation still indicates that PI is a reasonably reliable judge metric. Overall, these results suggest that judge scores track human perceptions of user-proxy quality.
4.4. Assistant Sensitivity and Rank Stability
LLM-as-judge scores can be confounded by the assistant model that the user proxy interacts with. To quantify this assistant sensitivity, we rerun evaluation on ChatbotArena while holding the judge fixed and varying the assistant among {GPT-4o, Claude-4-Sonnet, Gemini-2.5-Pro}. For each user-proxy model, we compute the three judge-based realism metrics (GTEval, PI, RNR) on the resulting rollouts.
Figure 5. Assistant sensitivity on ChatbotArena (scores). Mean $\pm$ 95% CI of judge-based realism metrics for each user-proxy when swapping the assistant model (color); the judge is fixed to Claude-4-Sonnet.
Figure 6. Rank stability under assistant swaps (ChatbotArena). Per-metric proxy ranks (1 = best) induced by the scores in Fig. 5 when varying the assistant model; stable lines indicate robust proxy ordering, while crossings reveal assistant-dependent rankings.
Figure 5 shows that absolute scores vary only slightly across assistants and are generally insufficient to change the overall ordering of user-proxy models. Figure 6 makes this explicit: ranks are stable (largely parallel trajectories) across assistants, with a single crossover in GTEval. Overall, assistant choice introduces mild variance in judge-based realism scores, but the resulting rankings remain stable most of the time.
4.5. Multi-Seed Robustness
LLM-as-judge metrics can fluctuate across repeated runs due to stochasticity in generation and evaluation. To assess the stability of our judge-based realism metrics, we perform a multi-seed analysis by repeating the benchmark five times with different random seeds for three different user-proxies while keeping the assistant and judge fixed.
Figure 7. Multi-seed robustness of judge-based realism metrics. For each user-proxy model, points show the per-seed mean score and error bars indicate the corresponding 95% CI on ChatbotArena; the assistant (gpt-4o) and judge (claude-4-sonnet) are fixed across runs.
[…]
4.6. Telemetry, Scaling, and Cost
@@
Based on our experiments, we further provide detailed analysis on the practical side of evaluation. Since most of the models we used are proprietary (incurs cost per million tokens), we provide statistical analysis on cost, latency and trade-off between cost vs performance. Per-sample telemetry (Fig. 8) shows that token usage is dominated by the judge, with the user-proxy & assistant contributing a smaller but non-negligible share; datasets differ markedly, OASST1 drives the highest token counts while ClariQ yields the largest end-to-end latency.
@@
Figure 8. Avg per-sample telemetry for GPT-4o as user-proxy & assistant, and judged by Claude-4-Sonnet across 4 datasets on 6 metrics (Fig. 2). Left: Token usage by role (darker: input, lighter: output). Right: Cumulative latency per episode.
@@
Figure 9. Cost-quality trade-off for PI: cost per evaluation (USD) vs. PI $\Delta w$ ($\uparrow$ better). Judge = Claude-4-Sonnet; assistant = GPT-4o; temperature = 0; cache off. Markers denote user-proxies; labels denote datasets. Dashed line: Pareto frontier. See Table 4 for model pricing.
@@
The cost-quality frontier (Fig. 9) clarifies trade-offs for PI evaluation: User-proxies Gemini-2.5-Pro and Claude-4-Sonnet offer attractive Pareto points (good PI $\Delta w$ at moderate cost), and GPT-5 generally incurs higher cost with weaker PI gain. Overall, user-proxy Gemini-2.5-Pro, along with Claude-4-Sonnet as judge, provides a balanced choice for large runs.
5. Conclusion & Future Work
MirrorBench provides a reproducible benchmark for evaluating the human-likeness of LLM-based user proxies, assessing how closely a proxy’s utterances match real human users across conversational tasks, independent of downstream task success. Across four public datasets, we evaluate five user-proxy LLMs using three judge-based realism metrics (GTEval, PI, RNR) together with human-anchored lexical diversity measures (MATTR, HD-D, Yule’s K). The resulting analyses highlight (i) systematic gaps between proxies and human users, (ii) a consistent realism-diversity trade-off that varies by dataset regime, and (iii) measurable sensitivity of absolute judge scores to the judge model, motivating calibration controls (HH/PP) and transparent multi-judge reporting. We also report token, latency, and cost profiles to make the evaluation trade-offs explicit when benchmarking at scale.
@@ 5. Conclusion & Future Work
Limitations and outlook. Judge-based metrics may reflect model-family bias and prompt sensitivity; while HH/PP controls provide useful context, they do not eliminate all sources of bias. Our experiments primarily use a fixed assistant configuration and limited randomization, and our coverage is restricted to four English-centric datasets. Finally, lexical diversity captures surface-level distributional variation and does not fully represent interaction-level diversity such as repair, initiative, or long-horizon consistency.
Overall, MirrorBench is intended as a benchmark for systematically comparing user proxies and for grounding progress on realistic user simulation in conversational AI.
Appendix A MirrorBench System
In this section, we explain the system aspects of the benchmarking framework. Although the metric definitions and execution may seem straightforward, substantial work went into creating a scalable, extensible system for flexible research use. We provide a detailed architectural blueprint in §A.1, explain the execution flow in §A.2, and provide metric implementation details in §A.3
A.1. Architecture
Figure 10. MirrorBench Architecture: Six-layer stack from low-level execution backends & persistence up through the core engine, plugin components, CLI & reporting, and task drivers. Top layers are user-facing; bottom layers are low-level infrastructure abstractions.
MirrorBench is a six-layer stack that cleanly separates infrastructure from evaluation logic (Fig. 10). Layers build upward, lower layers provide execution and data services; upper layers host pluggable components, task logic, and user interfaces.
A.1.1. Execution Backends & Persistence
@@
Execution Backend
It is responsible for taking decomposed units of tasks as input and send them for execution on the respective execution engine. The framework provides abstractions to support multiple execution backends to execute evaluation jobs at scale. The abstract interface provides uniform schema for implementations to extend the support for new execution engines. The framework supports fully-functional synchronous & asynchronous execution backends, and provides stub implementation to extend for Ray (Moritz et al., 2018) backend.
Persistence
All evaluation artifacts are persisted in a SQLite database organized around a three-level hierarchy: runs define experiment configurations, units specify individual evaluation tasks (parameterized by user-proxy, dataset, metric, and seed), and episodes capture per-trial results including execution duration, metric values, and telemetry. Aggregate statistics (mean, standard deviation, confidence intervals) are materialized in a separate metrics table to support efficient querying, while composite scores across multiple metrics are stored in scorecards for holistic comparisons. The complete schema is detailed in Table 2.
Table 2. Complete MirrorBench Run Database Schema
@@ A.1.2. Core Engine
Data Models
All the data structures for messages, episodes, evaluation units and run manifests are defined with typed data models using TypedDict and pydantic. We formally explain the context of these different data objects later.
Registry Builder
The framework provides certain high-level components which requires module-specific registries. This layer provides the abstract registry builder that supports creating metadata-aware registry for the downstream modules. Such registries can avoid silent errors since we can ensure certain constraints on the classes being registered with the metadata.
Configuration & Manifest Management
A typical evaluation job requires configuring user-proxy, datasets, metrics, scorecards etc. We define standard configuration schema for each of such components. Moreover, upon invocation of job or during dryrun, the system decomposes evaluation task into small evaluation units. In order to support parallelization (for async and remote backends), we persist self-contained information about the executable units as manifest files.
A.1.3. Orchestration